# On Selection of Cross-Section Averages in Non-stationary Environments

### v1

@@ Abstract
Information criteria (IC) are important tools in the literature of factor models that allow one to estimate a typically unknown number of latent factors. Although first proposed for the Principal Components setting in the seminal work by Bai and Ng, (2002), it has recently been shown that IC perform extremely well in Common Correlated Effects (CCE) and related setups with stationary factors. In particular, they can consistently select a sufficient set of cross-section averages (CAs) to approximate the factor space. Given that CAs can proxy nonstationary factors, it is tempting to believe that the consistency of IC continues to hold under such generality. This study is a cautionary tale for practitioners. We demonstrate formally and in simulations that IC has a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the CAs literature.
@@ 1 Introduction
Example 1. Common Correlated Effects (CCE). Let $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables and $\mathbf{y}_{i}$ is the dependent variable. Here, $K=k+1$, and $i$ indexes individuals, such as countries or households. Then (1.1) represents a classical CCE setup of Pesaran, (2006):
$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i},$ (1.2)
@@
where, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$. (1.1) nests (1.2) with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
Example 2. Distinct Correlated Effects (DCE). We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable:
$\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},,\quad\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ (1.3)
@@
as considered in Cui et al., (2022) or Stauskas and De Vos, (2025). Here, (1.1) nests (1.3) similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
@@
Example 3. Block Structure (FRED-MD-QD , EA-MD-QD). The set of regressors in Examples 1 and 2 can be high-dimensional, but follow a natural group structure with $k$ blocks and $m$ underlying factors. Let $i_{j}$ index series belonging to a block $j$ for $j=1,\ldots,k$. Then $\mathbf{Z}_{i}$ is $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$. Here, $N$ is the number of series per block. Blocks, such as monetary, industrial or fiscal series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets. These datasets have been very popular in factor-augmented forecasts (see e.g. Karabiyik et al.,, 2019, Stauskas and Westerlund,, 2022, or Morico and Stauskas,, 2025).
@@
In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). In Example 1, $\widehat{\mathbf{F}}$ is used as an additional regressor to merely control for the factors, and subsequently $\boldsymbol{\beta}$ is estimated with the least squares (LS), which constitutes the typical CCE estimator (see e.g. Karabiyik et al.,, 2017). In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025). In Example 3, coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g. Massacci and Kapetanios,, 2024).
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property holds as long as the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3). However, there are reasons to ensure that $m=K$. If $m<K$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In DCE, $\mathbf{F}_{\mathbf{x}}$ estimated with an excess error makes bootstrap inconsistent (see Stauskas and De Vos,, 2025). In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022). More generally, inclusion of too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis et al.,, 2021; Juodis,, 2022, for a discussion). In response to this, Margaritella and Westerlund, (2023) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be consistently applied in the pure CCE setting in Example 1. Subsequently, De Vos and Stauskas, (2024) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based selection (see also Ahn and Horenstein,, 2013). In that study, factors are stationary, as well. Another set of results concern fixed-$T$, where properties of $\mathbf{F}$ do not play an asymptotic role. Also, the method focuses solely on the number of factors, but cannot select the CAs. We leave analysis of this approach for the future research.
@@
In this study, we re-visit IC proposed in Margaritella and Westerlund, (2023) and De Vos and Stauskas, (2024). Given the power of CAs to proxy very general factor structures, it is only natural to evaluate IC under nonstationarity conditions. Contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity. We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC. Moreover, nonstationarity of $\mathbf{F}$ cannot be accommodated in the objective functions a priori, as their very design is based on the idea that properties of factors and idiosyncratics should be similar. This study can be seen as a caution to practitioners to transform data and not to extrapolate attractive properties of CAs should they need to choose a number thereof.
@@ 2 Econometric Setup
To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g=\mathrm{cols}(\mathbf{q}_{M})$ averages in practice. For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup. Moreover, let $\mathbf{q}_{M}=[\boldsymbol{0}_{g\times 1},\mathbf{q}_{M_{\mathbf{x}}}^{\prime}]^{\prime}$, then $\overline{\mathbf{Z}}\mathbf{q}_{M}=\overline{\mathbf{X}}\mathbf{q}_{M_{\mathbf{x}}}=\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ selects only from the set of $k$ regressors (or blocks), as required in Examples 2 and 3.222Throughout the remainder of the study, we use $m$ to denote the number of factors without the loss of generality. Clearly, this number can vary among our examples. Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rk}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rk}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$. IC under consideration are given by
@@
$p_{N,T}=\begin{cases}\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\\ \frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right).\end{cases}$ (2.3)
Here and further, MW represents Margaritella and Westerlund, (2023), while DVS stands for De Vos and Stauskas, (2024). $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs. Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$. Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002).
@@
      If $\tau=0$, then we let $\mathbf{e}_{i,t}=\mathbf{K}_{i}(L)\boldsymbol{\epsilon}_{i,t}=\sum_{j=0}^{\infty}\mathbf{K}_{i,j}\boldsymbol{\epsilon}_{i,t-j}$, where $\boldsymbol{\epsilon}_{i,t}$ is independent across $t$ with $\mathbb{E}(\boldsymbol{\epsilon}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\boldsymbol{\epsilon}_{i,t}\boldsymbol{\epsilon}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{\epsilon\epsilon,i,t}$ positive definite, $\mathbb{E}(\|\boldsymbol{\epsilon}_{i,t}\|^{4})<\infty$, and $\sum_{j=0}^{\infty}j^{1/2}\|\mathbf{K}_{i,j}\|<\infty$. Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size $-bd(b-d)$ with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite.
@@
Assumption 1 (a) parameterizes factors very flexibly and allows to perform comparative statics, since $\tau\to 1$ implies that the process becomes more persistent and approaches the usual local-to-unity process. We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allows us vary $\tau$ and document its effect on IC performance. This assumption also implies that $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$ which is a constant positive definite matrix. Its part (b) ensures that we return to the usual stationarity conditions under $\tau=0$. Assumption 2 (a) is split into two parts. If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to a zero matrix too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see also Lemma 3.1 in Magdalinos and Phillips,, 2009). For completeness, we present simulation evidence under serial correlation, as well, which does not demonstrate a major negative effect on IC. The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the factor loadings (rank condition in Assumption 4). The loadings are random as in e.g. Pesaran, (2006), but they can be treated as deterministic parameters, as well.
@@
Let us begin from the over-specification. In this case, the CAs are able to approximate all $m$ factors under Assumption 4. Also, for illustration purposes let $j=DVS$. Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator, which is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar argument applies to the objective function of MW. Thus, it is natural that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ is negligible. Moreover, the rate is already detected in Bai and Ng, (2002) under stationary $\mathbf{F}$ for PC, and it remains identical here. This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.3). Indeed, if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$, Proposition 1 is the key ingredient to show that $\mathbb{P}(\mathrm{IC}^{DVS}(M_{\mathbf{x}})-\mathrm{IC}^{DVS}(M_{\mathbf{x},0})>0)\to 1$, and the same statement holds for MW. This means that the probability that IC are minimized at some other set containing $M_{0}$ goes to 0. In general, under $M_{0}\subset M$, the usual CCE logic prevails deceptively.
@@
Under $M\subset M_{0}$, Assumption 4 is not satisfied, and the rank condition does not hold. This means that $m-g$ factors are not asymptotically projected out. Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a), but the objective function offers normalization by $T$ only. Exactly this imbalance induces divergence, and this implies that the probability that IC are minimized at some other set contained in $M_{0}$ does not necessarily go to 0. Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ (see the analysis of MW and DVS). Our Proposition 1 naturally provides the same result under $\tau=0$. Note that the failure of both IC is purely driven by the objective function and is not related to the penalty $p_{N,T}$. In fact, the only theoretical prescription for $p_{N,T}$ depends only on $C_{N,T}^{2}$ and not $\tau$, as given in (2.5).

### v2

@@ Abstract
Information criteria (IC) have been widely used in factor models to estimate an unknown number of latent factors. It has recently been shown that IC perform well in Common Correlated Effects (CCE) and related setups in selecting a set of cross-section averages (CAs) sufficient for the factor space under stationary factors. As CAs can proxy non-stationary factors, it is tempting to claim such generality of IC, too. We show formally and in simulations that IC have a severe underselection issue even under very mild forms of factor non-stationarity, which goes against the sentiment in the literature.
Keywords: Information criteria, factors, CCE, cross-section averages, non-stationary data
MOS subject classification: 91B84
@@ 1 Introduction
Example 1. Common Correlated Effects (CCE). Let $\mathbf{Z}_{i}=[\mathbf{y}_{i},\mathbf{X}_{i}]$, where $\mathbf{X}_{i}$ is a $T\times k$ matrix of explanatory variables and $\mathbf{y}_{i}$ is the dependent variable. Here, $K=k+1$, and $i$ indexes individuals, such as countries or households. Then (1.1) represents a classical CCE setup of Pesaran, (2006) with $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$ and $\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$. Here, $\boldsymbol{\beta}\in\mathbb{R}^{k}$ is the parameter vector, $\mathbf{V}_{i},\boldsymbol{\varepsilon}_{i}$ ($\boldsymbol{\Gamma}_{i}$, $\boldsymbol{\gamma}_{i}$) are the idiosyncratics (loadings). Both $\mathbf{X}_{i}$ and $\mathbf{y}_{i}$ load on $\mathbf{F}$, and this induces cross-section dependence among individuals and endogeneity as the error in $\mathbf{y}_{i}$ is correlated with $\mathbf{X}_{i}$. (1.1) nests this model with $\mathbf{C}_{i}=[\boldsymbol{\Gamma}_{i}\boldsymbol{\beta}+\boldsymbol{\gamma}_{i},\boldsymbol{\Gamma}_{i}]$ and $\mathbf{U}_{i}=[\mathbf{V}_{i}\boldsymbol{\beta}+\boldsymbol{\varepsilon}_{i},\mathbf{V}_{i}]$.
@@
Example 2. Distinct Correlated Effects (DCE). We can introduce distinct but correlated factors $\mathbf{F}=[\mathbf{F}_{\mathbf{x}},\mathbf{F}_{\mathbf{y}}]\in\mathbb{R}^{T\times(m_{x}+m_{y})}$ in Example 1, which makes the setup mode applicable: $\mathbf{y}_{i}=\mathbf{X}_{i}\boldsymbol{\beta}+\mathbf{F}_{\mathbf{y}}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$, $\mathbf{X}_{i}=\mathbf{F}_{\mathbf{x}}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}$ as considered in Cui et al., (2022) or Stauskas and De Vos, (2025). Here, (1.1) nests such model similarly to Example 1, but $\boldsymbol{\gamma}_{i}$ is modified accordingly, depending on whether $\mathbf{F}_{\mathbf{y}}\subset\mathbf{F}_{\mathbf{x}}$, vice versa or $\mathbf{F}_{\mathbf{x}}\cap\mathbf{F}_{\mathbf{y}}=\emptyset$.
@@
Example 3. Block Structure (FRED-MD-QD , EA-MD-QD). The regressors in Examples 1 and 2 can be high-dimensional, but follow a natural structure of $k$ blocks and $m$ underlying factors. Let $i_{j}$ index series in block $j$ for $j=1,\ldots,k$. Then $\mathbf{Z}_{i}=\mathbf{X}_{i}=\mathbf{F}\boldsymbol{\Gamma}_{i}+\mathbf{V}_{i}=[\mathbf{x}_{i_{1}},\dots,\mathbf{x}_{i_{k}}]^{\prime}$. Here, $N$ is the number of series per block. Monetary, industrial or fiscal blocks of series can be found in FRED-MD (see McCracken and Ng,, 2016), FRED-QD (see McCracken and Ng,, 2020) or EA-MD-QD (see Barigozzi et al.,, 2024) datasets. These datasets have been very popular in factor-augmented forecasts (see e.g. Karabiyik and Westerlund,, 2021, or Morico and Stauskas,, 2025).
@@
In all the cases above, we can estimate the factors by taking $K$ cross-section averages (CAs): $\widehat{\mathbf{F}}=\overline{\mathbf{Z}}=\mathbf{F}\overline{\mathbf{C}}+o_{p}(1)$, because CAs of the idiosyncratics are negligible under a wide range of empirically relevant assumptions (see e.g. Pesaran and Tosetti,, 2011). In Example 1, $\widehat{\mathbf{F}}$ is used as an additional regressor to merely control for the factors, and subsequently $\boldsymbol{\beta}$ is estimated with the least squares (LS), which constitutes the typical CCE estimator (see e.g. Karabiyik et al.,, 2017). In Example 2, only $\widehat{\mathbf{F}}_{\mathbf{x}}=\overline{\mathbf{X}}$ is used as a regressor, while $\mathbf{F}_{\mathbf{y}}$ is controlled by bootstrap (see Stauskas and De Vos,, 2025). In Example 3, the coefficient of $\widehat{\mathbf{F}}$ itself matters for forecast evaluation (see e.g. Massacci and Kapetanios,, 2024).
@@
A substantial advantage of CAs is their ability to estimate $\mathbf{F}$ irrespective of their time series properties. Indeed, the CCE estimator enjoys popularity, because it is consistent and asymptotically normal under very general $\mathbf{F}$ in large $N,T$ settings without any modifications (see e.g. Kapetanios et al.,, 2011, Westerlund,, 2018, Westerlund and Petrova,, 2018, or Stauskas,, 2023). This property is valid if the rank of $\overline{\mathbf{C}}$ is equal to $m$, which means that the number of factors cannot exceed the effective number of CAs ($K\in\{k,k+1\}$ in Examples 1-3). However, there are reasons to ensure that $m=K$. If $m<K$ and $TN^{-1}$ is bounded, CCE produces an asymptotic bias whose analytical correction is infeasible (see Karabiyik et al.,, 2017). This situation is common in macro datasets (see De Vos and Stauskas,, 2024). In DCE, $\mathbf{F}_{\mathbf{x}}$ estimated with an excess error makes bootstrap inconsistent (see Stauskas and De Vos,, 2025). In Example 3, popular tests of forecast accuracy comparison have highly non-standard asymptotic distributions (see Stauskas and Westerlund,, 2022). In general, too many CAs might even lead to a model that over-fits the data. In contrast, inclusion of too few generally leads to inconsistent estimates of the model parameters (see Juodis,, 2022, for a discussion). In response to this, Margaritella and Westerlund, (2023) (MW) were the first to demonstrate that Information Criteria (IC) of Bai and Ng, (2002) from Principal Components (PC) literature can be applied in the CCE setting in Example 1 under stationary factors to identify an optimal set of CAs. De Vos and Stauskas, (2024) (DVS) showed similar results applicable to more general Examples 2 and 3.111Juodis, (2022) considers eigenvalue-based methods of Ahn and Horenstein, (2013)) under stationarity. The focus there is the number of factors but not the optimal set of CAs. We leave analysis of this approach for the future research.
@@
MW stress that an assumption of stationary $\mathbf{F}$ is required only to simplify the proofs, hinting at a much greater generality of IC. The power of CAs to proxy general factor structures makes it natural to evaluate this statement and re-visit IC proposed in MW and DVS. We show that, contrary to the sentiment in the CCE/CAs literature, IC crumble under very mild forms of nostationarity. We demonstrate formally and in simulations that they exhibit a severe under-selection problem, which is independent of penalties used in the objective functions of IC. Moreover, nonstationarity of $\mathbf{F}$ cannot be accommodated in the objective functions a priori, as their very design is based on the idea that properties of factors and idiosyncratics should be similar. This study can be seen as a caution to practitioners to transform data and not to extrapolate attractive properties of CAs should they need to choose a number thereof.
@@ 2 Econometric Setup
To further operationalize our examples, we introduce $M$ which is a set of column indices of $\overline{\mathbf{Z}}$, and $\mathbf{q}_{M}\in\mathbb{R}^{K\times g}$ picks the corresponding $g$ averages in practice. For instance, $\overline{\mathbf{Z}}\mathbf{q}_{M}=\widehat{\mathbf{F}}_{M}$ selects $g$ out of $k+1$ CAs in the typical CCE setup. Moreover, let $\mathbf{q}_{M}=[\boldsymbol{0}_{g\times 1},\mathbf{q}_{M_{\mathbf{x}}}^{\prime}]^{\prime}$, then $\overline{\mathbf{Z}}\mathbf{q}_{M}=\overline{\mathbf{X}}\mathbf{q}_{M_{\mathbf{x}}}=\widehat{\mathbf{F}}_{M_{\mathbf{x}}}$ selects only from the set of $k$ regressors (or blocks), as required in Examples 2 and 3.222Throughout the remainder of the study, we use $m$ to denote the number of factors without the loss of generality. Clearly, this number can vary among our examples. Let accordingly $M_{0}$ ($M_{\mathbf{x},0}$) denote the set of averages from $\overline{\mathbf{Z}}$ ($\overline{\mathbf{X}}$) such that $\mathrm{rank}(\overline{\mathbf{C}}\mathbf{q}_{M_{0}})=\mathrm{rank}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x},0}})=m$, and $|M_{0}|=|M_{\mathbf{x},0}|=m$, where $|M|$ denotes cardinality of an arbitrary set $M$. IC under consideration are given by
@@
$p_{N,T}\in\left\{\frac{N+T}{NT}\ln\left(\frac{NT}{N+T}\right),\frac{N+T}{NT}\ln\left(C_{N,T}^{2}\right)\right\}$ (2.3)
where $\mathbf{M}_{\mathbf{A}}=\mathbf{I}-\mathbf{A}(\mathbf{A}^{\prime}\mathbf{A})^{+}\mathbf{A}^{\prime}$ is a projection matrix onto the orthogonal complement of $\mathbf{A}$, $\mathbf{A}^{+}$ is a Moore-Penrose (MP) inverse, while $\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ with $\widehat{\boldsymbol{\beta}}$ is obtained with the CCE estimator under all available $k+1$ CAs. Finally, $p_{N,T}$ is a penalty term which goes to 0 with $(N,T)\to\infty$. Examples of feasible and most popular penalties are given in (2.3) for $C_{N,T}=\mathrm{min}(\sqrt{N},\sqrt{T})$ (see more in Section 5 of Bai and Ng,, 2002). Note that (2.1) - (2.2) are functions of (a version of) the denominator of the CCE estimator, which is robust to general unknown factors (see Theorem 1 in Westerlund,, 2018).
@@
      If $\tau=0$, then we let $\mathbf{e}_{i,t}=\mathbf{K}_{i}(L)\boldsymbol{\epsilon}_{i,t}=\sum_{j=0}^{\infty}\mathbf{K}_{i,j}\boldsymbol{\epsilon}_{i,t-j}$, where $\boldsymbol{\epsilon}_{i,t}$ is independent across $t$ with $\mathbb{E}(\boldsymbol{\epsilon}_{i,t})=\mathbf{0}_{(k+1)\times 1}$, $\mathbb{E}(\boldsymbol{\epsilon}_{i,t}\boldsymbol{\epsilon}_{i,t}^{\prime})=\boldsymbol{\Sigma}_{\epsilon\epsilon,i,t}$ positive definite, $\mathbb{E}(\|\boldsymbol{\epsilon}_{i,t}\|^{4})<\infty$, and $\sum_{j=0}^{\infty}j^{1/2}\|\mathbf{K}_{i,j}\|<\infty$. Also, $\boldsymbol{\xi}_{t}=\mathrm{vec}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}-\mathbb{E}\left(N\overline{\mathbf{e}}_{t}\overline{\mathbf{e}}_{t}^{\prime}\right)\right)$ is strong mixing with coefficients of size $-bd(b-d)$ with $b>4$ and $b>d>2$, $\mathbb{E}(\|\boldsymbol{\xi}_{t}\|^{b})<\infty$, and $\lim_{T\to\infty}T^{-1}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}(\boldsymbol{\xi}_{t}\boldsymbol{\xi}_{s}^{\prime})$ is positive definite. Here and throughout, $\left\|\mathbf{A}\right\|=\sqrt{\mathrm{trace}(\mathbf{A}^{\prime}\mathbf{A})}$ is Frobenius norm.
@@
Assumption 1 (a) treats factors very flexibly and allows us perform comparative statics, as $\tau\to 1$ increases persistence of the process. We are not interested in a specific model for $\mathbf{F}$, but Magdalinos and Phillips, (2009) allow us vary $\tau$ and give theoretical guarantees, such as $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{F}\to_{p}\boldsymbol{\Sigma}_{\mathbf{F}}$, which is a constant positive definite matrix. Part (b) ensures that we return to the usual stationarity conditions under $\tau=0$. Assumption 2 (a) is split into two parts. If the idiosyncratics $\{\mathbf{e}_{i,t}\}$ are correlated over time, for $\mathbf{E}_{i}=[\mathbf{e}_{i,1},\ldots,\mathbf{e}_{i,T}]^{\prime}$, $\frac{1}{T^{1+\tau}}\mathbf{F}^{\prime}\mathbf{E}_{i}$ converges to zero too slow for our purposes, which would dominate and obscure the very effect of factor persistence on IC, and we would like to avoid this (see Lemma 3.1 in Magdalinos and Phillips,, 2009). We present simulations under serial correlation in the Supplement, which does not demonstrate a major negative effect on IC. The rest of the assumptions are standard, as they ensure mutual independence of factors and the error terms (Assumption 3), weak cross-section dependence of idiosyncratics (Assumption 2 (b)) and informativeness of the loadings (rank condition in Assumption 4). The loadings are random as in e.g. Pesaran, (2006), but they can also be treated as deterministic parameters.
@@
Let us begin from the over-specification. In this case, the CAs are able to approximate all $m$ factors under Assumption 4. Also, for illustration purposes let $j=DVS$. Then $\mathcal{V}_{DVS}$ (and $\mathcal{V}_{DVS}^{0})$ is exactly (log-determinant of) the denominator of the CCE estimator. It uses the fact that under $M_{0}\subset M$ we have $\mathbf{F}=\overline{\mathbf{X}}(\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{\mathbf{x}}})^{+}+o_{p}(1)$, and by inserting this into $\mathbf{X}_{i}$, we can show that
$\mathcal{V}_{DVS}=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{M}_{\mathbf{F}}\mathbf{V}_{i}\right)+o_{p}(1)=\ln\mathrm{det}\left(\frac{1}{NT}\sum_{i=1}^{N}\mathbf{V}_{i}^{\prime}\mathbf{V}_{i}\right)+O_{p}(T^{-1}),$ (2.6)
@@
whose argument is consistent for a positive definite matrix under general factors (see Westerlund,, 2018), and a similar logic applies to the objective function of MW. Thus, it is natural that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}$ is negligible. Moreover, the rate is already detected in Bai and Ng, (2002) under stationary $\mathbf{F}$ for PC, and it remains identical here. This rate is instructive and determines the properties of penalty $p_{N,T}$, which is the key motivation behind the functional forms in (2.3). Indeed, if $p_{N,T}\to 0$ and $p_{N,T}C_{N,T}^{2}\to\infty$ as $(N,T)\to\infty$, Proposition 1 is the key ingredient to show that $\mathbb{P}(\mathrm{IC}^{DVS}(M_{\mathbf{x}})-\mathrm{IC}^{DVS}(M_{\mathbf{x},0})>0)\to 1$, and the same statement holds for MW. This means that the probability that IC are minimized at some other set containing $M_{0}$ goes to 0. In general, under $M_{0}\subset M$, the usual CCE logic prevails deceptively.
@@
Under $M\subset M_{0}$, the rank condition in Assumption 4 is not satisfied. This means that $m-g$ factors are not asymptotically projected out, and the representation in (2.6) is invalid. Given that $\mathcal{V}_{j}$ is a quadratic form for both $j\in\{MW,DVS\}$, their second empirical moment is of the order $O_{p}(T^{1+\tau})$ as implied by Assumption 1 (a). However, the objective function offers normalization by $T$ only. Exactly this imbalance induces divergence, and this implies that the probability that IC are minimized at some other set contained in $M_{0}$ does not necessarily go to 0. Under stationary $\mathbf{F}$, however, it is clear that $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}=O_{p}(1)$ and, in fact, $\mathcal{V}_{j}-\mathcal{V}_{j}^{0}\to_{p}c>0$ as $(N,T)\to\infty$ as required (see the analysis of MW, DVS and Bai and Ng,, 2002). Our Proposition 1 naturally provides the same result under $\tau=0$. Note that the failure of both IC is purely driven by the objective function, and it complicates the role of $p_{N,T}$. If $M_{0}\subset M$, $p_{N,T}$ should simply vanish slower than $C_{N,T}^{-2}$, but under $M\subset M_{0}$ the penalty has to counter divergence.